# Supplement to “Steering LLMs Towards Unbiased Responses: A Causality-Guided Debiasing Framework”

### v1

@@ Abstract
Large language models (LLMs) can easily generate biased and discriminative responses. As LLMs tap into consequential decision-making (e.g., hiring and healthcare), it is of crucial importance to develop strategies to mitigate these biases. This paper focuses on social bias, tackling the association between demographic information and LLM outputs. We propose a causality-guided debiasing framework that utilizes causal understandings of (1) the data-generating process of the training corpus fed to LLMs, and (2) the internal reasoning process of LLM inference, to guide the design of prompts for debiasing LLM outputs through selection mechanisms. Our framework unifies existing de-biasing prompting approaches such as inhibitive instructions and in-context contrastive examples, and sheds light on new ways of debiasing by encouraging bias-free reasoning. Our strong empirical performance on real-world datasets demonstrates that our framework provides principled guidelines on debiasing LLM outputs even with only the black-box access.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs), which are trained on massive text corpora, have been found to exhibit concerning levels of social biases (Sheng et al., 2019; Gonen & Goldberg, 2019; Schick et al., 2021; Bender et al., 2021; Dodge et al., 2021). The unchecked biases can potentially perpetuate and amplify societal inequities, leading to unfair or even unethical outcomes. This issue is particularly significant as LLMs become more capable and start to serve as foundational components in decision-making systems across various sectors such as healthcare and education. Many debiasing approaches have been proposed to tackle this issue, for instance, direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023). For various reasons such as security and business interests, many of the most capable LLMs are closed-sourced, for instance, GPT-4 (Achiam et al., 2023), Gemini (Anil et al., 2023), Claude 2 (Anthropic, 2023), where the general public do not have access to models’ internal structures or parameters. Thus, prompting-based techniques largely become the only viable option to mitigate bias when using closed-sourced LLMs.
Figure 1: Different reasoning to answer a coreference question. A biased answer may be due to the use of a gender shortcut, while a bias-free answer is made by considering proper world knowledge given the circumstances. [^2] The use of demographic information does not necessarily indicate the reasoning is biased: sometimes certain demographic information (e.g., gender) should be considered for situations such as making medical decisions.
@@
In this work, we focus on prompting techniques to steer LLMs towards unbiased responses. To study this, we notice that obtaining unbiased responses essentially boils down to the process of selecting proper pieces from the model’s internal representations and knowledge to make the decision. Let us consider a simple task of resolving the coreference of a gender pronoun in a given sentence. A model may output a biased answer by using a gender shortcut potentially learned from the imbalanced representations in the training data. For example, in Figure 2(a), the model may associate the pronoun “he” with “physician” rather than “secretary” due to social biases reflected in the training data – the gender distributions are disproportional for different occupations. We denote such reasoning process as biased reasoning, where the model selects internal representations in an improper way as this specific demographic information (gender) is unrelated to solving the coreference task. We provide more detailed discussions on selection mechanisms in Section 3.
Figure 2: Different prompting-based methods guided by one or both intuitions: discouraging ([-]) biased reasoning, and encouraging ([+]) bias-free reasoning.
@@
Most of the existing prompting-based debiasing methods focus on discouraging biased reasoning, for instance, using explicit prompts to avoid using biased associations or prohibiting the utilization of gender information when making the decision. While these approaches help to a certain extent, another overlooked prompting-based strategy is to encourage bias-free reasoning by only selecting proper pieces related to the nature of the question. For example, Figure 2(b) demonstrates one possible way of conducting bias-free reasoning by comparing the likelihood of two situations occurring in real life and drawing on the more plausible situation to infer the coreference resolution.
We reveal the essential role of selection mechanisms in the interplay between LLM’s internal reasoning process and different strategies for designing the external prompts. We first construct a detailed causal model on the data generation process of the training corpus to identify ways biases may be smuggled in the pretraining phase. Then we construct a causal model of the LLM’s potential reasoning process and connect the two causal models by analyzing how different input prompts could modulate the LLM’s output through selection mechanisms. Building upon the above causal understandings, we introduce a causality-guided framework to debias LLMs, underpinned by the two intuitions: (1) discouraging biased reasoning  and (2) encouraging bias-free reasoning. Current prompting-based debiasing methods can also be viewed as instantiating one or both of the intuitions, as illustrated in Figure 2.
@@
Using the causality-guided debiasing framework, we propose prompt-design strategies and further conduct a variety of empirical studies to analyze their effectiveness. We find that prompts employing both intuitions significantly outperform existing prompting-based debiasing methods. The strong empirical results clearly demonstrate the effectiveness of the causality-guided debiasing framework, even when we only have black-box accesses.
Our contributions can be summarized as follows:
  • We construct detailed causal modelings for both the underlying data-generating process of the training corpus and the LLM reasoning process. We note that selection mechanisms play an essential role in identifying how the LLM’s output could be modulated by different prompts.
  • We formulate a causality-guided debiasing framework. Through causal understandings of data-generating processes, we identify conditions that strategies for designing prompts should satisfy to debias in a principled and effective way. The framework also unifies existing prompting-based debias methods.
  • Using these strategies, we show strong empirical results on debiasing various social biases across different LLMs, demonstrating the clear benefit of our framework.
@@ 2 Preliminaries
In this section, we provide a brief introduction to causal modeling and causal reasoning (Section 2.1). We also provide a motivating example to illustrate how the selection mechanism can reshape dependence patterns within the different data (Section 2.2). [^3] We discuss related works in detail in Appendix A.
@@ 2.1 A Brief Introduction to Causality
For two random variables $X$ and $Y$, $X$ is a cause of $Y$ if there is a change in the distribution of $Y$ when we apply an intervention on $X$ while holding all other variables fixed (Spirtes et al., 1993; Pearl, 2009). We can represent causal relations among variables with a directed acyclic graph (DAG), where nodes represent variables, and edges represent direct causal relations between variables. We denote the direct causal relation between the ordered pair $(X,Y)$ by a directed edge $X\rightarrow Y$.
Local causal modules in a DAG, which characterize the causal relations between the corresponding variable and its direct causes, do not interfere with each other because of causal modularity. This property is also known as the exogeneity (Engle et al., 1983), or the independence of causal mechanism (Peters et al., 2017), resulting directly from the causal Markov condition for the DAG (Spirtes et al., 1993; Pearl, 2009). [^4] There are additional classes of graphs considered in causality literature, for example, directed cyclic graphs (DCGs) (Spirtes, 1995), ancestral graphs (Richardson & Spirtes, 2002), and so on. In this paper, we consider causal processes that can be modeled by a DAG. Other graph classes are beyond the scope of our work. In the context of language processing, the definition of a variable representing text or tokens is relatively abstract compared to the statistical notion of a random variable in tabular data. Within the scope of this work, we use the terms “variable” and “node” interchangeably when the context permits clear understandings.
2.2 A Motivating Example of Selection Mechanisms
Ideally, we would like samples to be drawn uniformly from the underlying population of interest. However, in practice, it is very common that the probability of including certain data points in the corpus depends on the characteristics of the data points themselves.
Previous literature has investigated selection mechanisms from different perspectives, for instance, the influence of selection bias on statistical inference in economic and sociological studies (Heckman, 1979, 1990; Winship & Mare, 1992), causal discovery when there are selection variables and latent common causes (Spirtes et al., 1995; Zhang, 2008), the identification and estimation of functional causal models when selection exists (Zhang et al., 2016), the identifiability of causal effect in the presence of selection bias from the graphical condition perspective (Bareinboim & Pearl, 2012; Bareinboim & Tian, 2015; Correa et al., 2019) and from the potential outcome perspective (Hernán & Robins, 2020), and the identification of the existence of selection bias from observational data under certain functional assumptions (Kaltenpoth & Vreeken, 2023).
@@ 2.2 A Motivating Example of Selection Mechanisms
Let us consider an example in the context of medical care. As illustrated in Figure 3, observed variables $(X_{1},X_{2})$ denote diseases, and $(Y_{1},Y_{2})$ denote corresponding symptoms. Apart from them, there exist potential binary selection variables $S_{i}$’s ($i\in\{1,2,\ldots,5\}$), where $S_{i}=1$ denotes being selected. $X_{1}$ ($X_{2}$) is the direct and only cause of $Y_{1}$ ($Y_{2}$), and the two diseases are unrelated in the general population, i.e., when none of the $S_{i}$’s exists. We use solid edges to represent causal relations among observed variables and dashed edges for those pertaining to selection mechanisms.
For instance, $S_{2}$ is an example of outcome-dependent selection (Zhang et al., 2016), which can be selecting individuals with symptom $Y_{1}$ from the general population. The conditional probability $P(Y_{1}\mid X_{1},S_{2}=1)$ in the selected data typically differs from its counterpart $P(Y_{1}\mid X_{1})$ in general population (Zhang et al., 2016). As another example, $S_{4}$ can denote the selection mechanism of only considering the hospital in-patient data. This signifies the setting of the well-known Berkson’s Paradox (Berkson, 1946), where two unrelated diseases $(X_{1},X_{2})$ appear to be correlated in the hospital data, simply because the data only contains selected patients who have at least one symptoms in $(Y_{1},Y_{2})$.
Figure 3 shows that the selection can be based solely on the cause (e.g., $S_{1}$), solely on the effect (e.g., $S_{2}$), or on both (e.g., $S_{3}$). Meanwhile, the selection can be based on variables that are not causally related in the general population (e.g., $S_{4}$ and $S_{5}$). Selection mechanisms can reshape dependence patterns among involved variables, and as a consequence, potentially change downstream outputs as well. We will see in Section 3 that such property of selection also applies to natural language processing (NLP) contexts.
@@
(b) LLM reasoning process given prompt
(c) LLM reasoning process (annotated copy)
Figure 4: Causal graphs with respect to different data-generating processes. Panel (a) presents the underlying data-generating process of the training data corpus. Panels (b) – (c) present the reasoning process of the LLM. Since the conditions specified by the prompt are always conditioned upon when LLM generates the output, the internal reasoning is modulated by the external prompt. In panel (c) we include annotations to highlight causal pathways along which the information flow from demographic representation to LLM potential output is not regulated.
3 Causality-Guided Debiasing Framework
We propose our debiasing framework guided by causal understandings of involved data-generating processes. In Section 3.1, we present detailed causal models of both the underlying data-generating process for the training corpus and the LLMs’ reasoning process. We reveal that the LLM’s reasoning process is essentially an interplay between its internal representations and conditions specified by the external prompt designs, in which selection mechanisms play a key role. In Section 3.2, we identify different sets of conditions for prompt designs, which lead to prompting strategies that are not only intuitive but also theoretically grounded.
3.1 Detailed Causal Modeling of Processes
We present the causal modeling of the underlying data-generating process of the training data corpus (Section 3.1.1) and the reasoning process of LLMs (Section 3.1.2).
3.1.1 Underlying Data Generating Process of Text in Training Data Corpus
@@
The training data corpus often reflects historical discrimination, and selection mechanisms are commonly involved. For instance, gender stereotype in occupations arises not due to the existence of a direct causal relation or a common cause between gender and occupation, but due to an underlying selection mechanism. Specifically, in the training data, among all possible combinations between gender (e.g., male and female) and occupation (e.g., CEO and secretary), there is a tendency to associate CEO more often with male, and secretary with female. This phenomenon occurs because the training data is a subset selected from an imaginary data corpus that is ideally diverse and comprehensive.
@@
More generally, in Figure 4(a) we model causal relations in the underlying data-generating process of training data corpus (e.g., text scraped from the internet). Other than demographic information (e.g., race, gender, age), the causal graph contains additional variables of interest. We use “scenario” to represent the practical situation or context that sets the background of the text (e.g., medical care, hiring). We use “entity” to denote participants or stakeholders involved in the scenario, for instance, a patient may be an entity in the medical care scenario and a secretary may be an entity in the hiring scenario. The selection variable $S$ explicitly models the association between demographic information and entity, which is recognized as a major type of stereotype in NLP (Sweeney, 2013; Bolukbasi et al., 2016; Zhao et al., 2018; Tamkin et al., 2023). There are different types of texts in Figure 4(a). “Demographic-agnostic fact” denotes the text that does not explicitly contain demographic information. “Derivational text based on fact” denotes the text derived from fact, e.g., inference according to definition, fact-check Q&A, and restatement without altering factual contents. “Demographic-aware text” denotes the text where demographic information appears explicitly.
3.1.2 Reasoning Process of LLMs
Section 3.1.1 introduces causal modeling of the underlying data-generating process to characterize how discrimination is instantiated in LLMs’ training data corpus. The goal and expectation of the training is for LLM models to capture dependence patterns in the data corpus, and to extract and internalize the related knowledge. Therefore, we assume that the internal reasoning process of LLMs shares similarities with the underlying data-generating process of data corpus. Under this mild assumption, we formulate the reasoning process of how LLMs generate outputs based on input prompts in Figure 4(b). In particular, we reveal the interplay between internal representations and conditions specified by external inputs, and more importantly, how LLM outputs are shaped and modulated by different prompt designs through selection mechanisms.
In Figure 4(b), we use dotted contours to distinguish LLMs’ internal representations from the actual external information in the data corpus, e.g., the contrast between “demographic representation” in Figure 4(b) and “demographic information” in Figure 4(a). Notice that the internal nodes are not directly observable or accessible. We use double-stroke contours to indicate selection mechanisms, directed edges to represent direct causal relations, and dashed edges with hollow arrowheads to denote selection mechanisms.
A “prompt” serves as an external input to LLMs but not as a direct cause of internal representations. This is because the internal knowledge and representations exist beforehand, making them irrelevant to whether a specific prompt is provided. The prompt also does not act as an indicator for causal interventions. Because internal nodes are not directly observable or accessible, one cannot set them to certain values via hard interventions (Spirtes et al., 1993; Pearl, 2009; Peters et al., 2017; Hernán & Robins, 2020), or change the functional behavior of causal modules via soft interventions (Eberhardt & Scheines, 2007; Huang et al., 2020; Correa & Bareinboim, 2020). However, the prompt directly changes the selection variable “prompt properly considered” (PPC) through its designs. If the LLM model is well-trained and well-aligned, we can expect that PPC is always conditioned upon when the model produces the output, and the “LLM potential output” will become the actual output we obtain from the model. Just like the motivating example where selection can reshape dependence patterns among involved variables (Section 2.2), the prompt can influence the LLM reasoning process in significant ways by specifying conditions in selection mechanisms.
3.1.3 A Remark on the Modeled Processes
Figure 4(a) and Figure 4(b) each have their generating process of interest and distinct emphases. Although not pertaining to LLM reasoning itself, the underlying generating process modeled in Figure 4(a) provides hints on local causal modules of interest that prompt designs can specifically attend to for debiasing purposes. When we consider these hints with reference to the LLM reasoning process modeled in Figure 4(b), we can identify debiasing strategies that are both intuitive and theoretically grounded, as we will see in more detail in Section 3.2. The two detailed causal models complement each other, both of which are essential for understanding the source of bias, and furthermore, effectively debiasing LLMs.
3.2 Debiasing Guided by Causal Understandings
We present our prompt-based LLM debiasing framework guided by causal understandings of the related data-generating processes. Our core idea is to formulate conditions that should be specified in the prompt design, such that through the influence of selection mechanisms in the LLM reasoning process, one can effectively debias LLM outputs.

### v2

# Prompting Fairness: Integrating Causality to Debias Large Language Models

@@ Abstract
Large language models (LLMs), despite their remarkable capabilities, are susceptible to generating biased and discriminatory responses. As LLMs increasingly influence high-stakes decision-making (e.g., hiring and healthcare), mitigating these biases becomes critical. In this work, we propose a causality-guided debiasing framework to tackle social biases, aiming to reduce the objectionable dependence between LLMs’ decisions and the social information in the input. Our framework introduces a novel perspective to identify how social information can affect an LLM’s decision through different causal pathways. Leveraging these causal insights, we outline principled prompting strategies that regulate these pathways through selection mechanisms. This framework not only unifies existing prompting-based debiasing techniques, but also opens up new directions for reducing bias by encouraging the model to prioritize fact-based reasoning over reliance on biased social cues. We validate our framework through extensive experiments on real-world datasets across multiple domains, demonstrating its effectiveness in debiasing LLM decisions, even with only black-box access to the model.
Prompting Fairness: Integrating Causality to Debias Large Language Models
11footnotetext: Equal contribution. Correspondence to: Jingling Li <jinglingli1024@gmail.com> and Zeyu Tang <zeyutang@cmu.edu>. The work was done while JL and YL were working at ByteDance Research.
@@ 1 Introduction
Large language models (LLMs) trained on massive text corpora have been found to exhibit concerning levels of social biases (Sheng et al., 2019; Gonen & Goldberg, 2019; Schick et al., 2021; Bender et al., 2021; Dodge et al., 2021). The unchecked biases can potentially perpetuate and amplify societal inequities, leading to unfair or even unethical outcomes. This issue is particularly significant as LLMs become more capable and start to serve as foundational components in decision-making systems across various sectors such as healthcare and education. Many debiasing approaches have been proposed to tackle this issue, for instance, direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023). For various reasons such as security and business interests, many of the most capable LLMs are closed-source models, for instance, GPT-4 (OpenAI, 2023), Gemini (Anil et al., 2023), Claude (Anthropic, 2024), where the general public do not have access to models’ internal structures or parameters. Therefore, prompting-based techniques largely become the only viable option to mitigate bias when using closed-source LLMs.
@@
In this work, we focus on prompting techniques that mitigate bias in LLMs’ decisions. We observe that unbiased decision-making in an LLM essentially hinges on the selection and utilization of appropriate dependence patterns within its internal representations or knowledge. Let us consider a simple coreference resolution task on gender pronouns in a given sentence. A model may output a biased answer via a gender shortcut potentially learned from the unbalanced training data. For example, in Figure 1(a), the model may associate the pronoun “she” with “teacher” rather than “driver” due to social biases reflected in the training data–the gender distributions are disproportional for different occupations. We denote such reasoning process as biased reasoning, where the model selects and utilizes inappropriate dependence patterns involving specific social information (gender in this example) that should actually be irrelevant to solving the coreference task.
(a) Different reasoning to a coreference question
(b) A systematic way to promote fact-based reasoning
Figure 1: Panel (a): A biased answer may be due to the use of a gender shortcut, while a fact-based answer is made by considering proper world knowledge given the circumstances. Panel (b): we describe how to systematically generate prompts that encourage fact-based reasoning. We would like to note that using social category information does not necessarily indicate that the reasoning is biased: sometimes, certain social category information should be considered, e.g., gender in medical treatments. We call such dependence neutral dependence, while in this work, we focus on objectionable/problematic ones (Tang et al., 2024).
@@
Most existing prompting-based debiasing methods focus on discouraging biased reasoning, for instance, using explicit instructions to avoid the usage of biased association when making the decision. While these approaches help to a certain extent, another overlooked debiasing strategy is to encourage fact-based reasoning by only selecting proper dependence patterns related to the nature of the question. For example, in Figure 1(a), we ask the model to first compare the likelihood of two situations occurring in real life and draw on the more likely situation to infer the coreference resolution. In Figure 1(b), we also provide a systematic approach to encourage fact-based reasoning.
We formalize the above intuitions into a causality-guided debiasing framework, offering a novel perspective on how social information influences an LLM’s decision-making through different causal pathways. Within this proposed framework, we demonstrate how selection mechanisms can introduce bias into the training data while also help mitigate bias in LLMs’ decisions. Specifically, we present principled prompting-based strategies that regulate different causal pathways via employing selection mechanisms, and these strategies can be applied to (1) discourage biased reasoning, and (2) encourage fact-based reasoning. Current prompting-based debiasing methods can be viewed as instantiating one or both of these goals.
@@
We conduct extensive empirical studies on real-world datasets across multiple domains, and our proposed strategies significantly outperform existing prompting-based debiasing methods. The strong empirical results clearly demonstrate the effectiveness of our causality-guided debiasing framework, even when we only have black-box accesses. Our contributions can be summarized as follows:
  • We develop a causality-guided debiasing framework that provides a novel perspective on how social information can influence LLMs’ decisions through different causal pathways.
  • Using the proposed framework, we present principled prompting-based debiasing strategies that utilize selection mechanisms to regulate the influence of bias. These strategies aim to discourage biased reasoning and encourage fact-based reasoning, offering a unified view of existing prompting-based methods and addressing the gap in promoting fact-based reasoning.
  • We conduct extensive empirical studies and demonstrate the effectiveness of our framework in debiasing LLMs’ decisions across various social categories, highlighting both its theoretical foundations and practical significance in debiasing LLMs’ decisions.
@@ 2 Preliminaries
In this section, we introduce our problem setting, defining what we mean by an unbiased decision for large language models (Section 2.1). Then we provide a brief introduction to causal modeling and causal reasoning (Section 2.2). We also provide a motivating example to illustrate how the selection mechanism can reshape dependence patterns among different variables (Section 2.3).
2.1 Problem Settings
[…]
@@ 2.2 A Brief Introduction to Causality
For two random variables $X$ and $Y$, $X$ is a direct cause of $Y$ if there is a change in the distribution of $Y$ when we apply an intervention on $X$ while holding all other variables fixed (Spirtes et al., 1993; Pearl, 2009). We can represent causal relations among variables with a directed acyclic graph (DAG), where nodes represent variables, and edges represent direct causal relations between variables. We denote the direct causal relation between the ordered pair $(X,Y)$ by a directed edge $X\rightarrow Y$. In the context of language processing, the definition of a variable representing text or tokens is relatively abstract compared to the statistical notion of a random variable in tabular data. Within the scope of this work, we use the terms “variable” and “node” interchangeably when the context permits clear understandings.
2.3 Remarks on Selection Mechanisms: An Illustrative Example
Ideally, we would like samples to be drawn uniformly from the underlying population of interest. However, in practice, the probability of including certain data points in the training corpus often depends on their characteristics where selection mechanisms are commonly involved. [^1] A thorough literature review on selection mechanisms can be found in Appendix A.1.
@@ 2.3 Remarks on Selection Mechanisms: An Illustrative Example
Consider an example in the context of medical care: as illustrated in Figure 2, the observed variables $(X_{1},X_{2})$ denote diseases and $(Y_{1},Y_{2})$ represent corresponding symptoms. There can exist potential binary selection variables $S_{i}$’s ($i\in\{1,2,\ldots,5\}$), where $S_{i}=1$ indicates the occurrence of that selection. $X_{1}$ ($X_{2}$) is the direct and only cause of $Y_{1}$ ($Y_{2}$), and the two diseases are uncorrelated in the general population, i.e., when none of the $S_{i}$’s exists. Solid edges in the figure denote causal relations among observed variables, while dashed edges represent the ones pertaining to selection mechanisms.
For instance, $S_{2}$ is an example of outcome-dependent selection (Zhang et al., 2016), which can be selecting individuals with symptom $Y_{1}$ from the general population. The conditional probability $P(Y_{1}\mid X_{1},S_{2}=1)$ in the selected data typically differs from its counterpart $P(Y_{1}\mid X_{1})$ in the general population (Zhang et al., 2016). As another example, $S_{4}$ demonstrates the selection solely based on the hospital in-patient data, a setting of the well-known Berkson’s Paradox (Berkson, 1946), where two unrelated diseases $(X_{1},X_{2})$ appear to be correlated in the hospital data, simply because the sample includes only admitted patients who have at least one symptoms in $(Y_{1},Y_{2})$.
[…]
3 Causality-Guided Debiasing Framework
This section presents our causality-guided debiasing framework. Specifically, we provide a causal modeling on the data generating process of LLMs’ training corpora in Section 3.1, highlighting how bias can be introduced into the training data by selection mechanisms. In Section 3.2, we demonstrate how social category information can impact LLMs’ decisions via different causal pathways. Section 3.3 further reveals how selection mechanisms can regulate the information flow over these pathways by reshaping dependence patterns. The regulations over different causal pathways leads to prompting strategies with distinct objectives aimed at effectively mitigating biases in LLMs. Moreover, we prove that when these objectives are satisfied, we can completely remove the influence of social category information from the LLMs’ decisions, which allows the LLMs’ decision to be conditionally independent of the social information, given the presence of appropriate selection mechanisms.
@@ 3 Causality-Guided Debiasing Framework
(b) How input prompts modulate LLM decisions
Figure 3: Causal graphs for different data generating processes. We use double-stroke contours to indicate selection variables, solid edges to represent causal relations among observed variables or internal representations, and dashed edges for those pertaining to selection mechanisms. Panel (a) presents the underlying data generating process of the training data corpus. Panel (b) presents a causal perspective on how LLM’s decisions are related to internal representations and modulated by external prompts. We highlight in light coral to denote pathways along which social category information can influence the LLM’s decision.
3.1 Underlying Data Generating Process of LLMs’ Training Corpora
@@
The training data of LLMs often reflects historical discrimination or stereotypes, and selection mechanisms are commonly involved. For instance, gender stereotype in occupations arises not due to the existence of a direct causal relation or a common cause between gender and occupation, but rather from an underlying selection mechanism: in the training data, among all possible combinations between gender (e.g., male and female) and occupation (e.g., CEO and secretary), there is a tendency to associate CEO more often with male, and secretary with female due to historical or societal biases. This phenomenon may occur when the training data is a subset selected from an imaginary data corpora that is ideally diverse and comprehensive, where there is no historical discrimination or stereotype when people generate text content. Therefore, the selection mechanism between gender and occupations introduces bias into the training data, which was fed to LLMs during the pretraining stage.
@@
More generally, in Figure 3(a), we provide a causal graph to denote the underlying data generating process of training corpora (e.g., text scraped from the internet). The causal graph contains additional variables of interest besides information related to a social category (e.g., age. gender, and race): “scenario” represents the practical situation or context where a decision is needed (e.g., medical care, hiring); “entity” denotes the participants or stakeholders involved in the scenario (e.g., a patient in the medical care scenario, or a secretary in the hiring scenario); and the selection variable $S$ explicitly models the associations of social categories with particular entities and scenarios, which are recognized as a major type of stereotypes in NLP (Sweeney, 2013; Bolukbasi et al., 2016; Zhao et al., 2018; Tamkin et al., 2023). There are also different types of texts in Figure 3(a): “social-agnostic fact” denotes the text that does not explicitly contain information related to a social category of interest (e.g., base question in Figure 1(b)); “social-salient text” denotes the text where information related to the social category appears explicitly (e.g., original question in Figure 1(b)); and “derivational text based on fact” denotes the text or decision derived from factual evidence (e.g., inference according to fact-based reasoning).
3.2 How Social Category Information Can Impact LLM’s Decisions